OpenStudy (anonymous):

Suppose R is the region in the first quadrant bounded by the curves y=3/x y=(x^2)/9 the volume of the solid resulting when revolving R about the y−axis.

OpenStudy (anonymous):

You wanna graph the curves first?

OpenStudy (anonymous):

ok so i feel like i can confidently do it if it we're around the Y axis, but that x=1 is what is throwing me off

OpenStudy (anonymous):

What do you mean \(x=1\)?

OpenStudy (anonymous):

lol oh wow, i never typed that part in

OpenStudy (anonymous):

Revised :Suppose R is the region in the first quadrant bounded by the curves y=3/x y=(x^2)/9 and x=1 . Find the volume of the solid resulting when revolving R about the y−axis.

OpenStudy (anonymous):

Let's graph first.

OpenStudy (anonymous):

I would let \[ \frac 3x=\frac{x^2}{9} \]

OpenStudy (anonymous):

Simplifying that seems to get \(x=3\), when they intersect.

OpenStudy (anonymous):

yeah i have the same thing

OpenStudy (anonymous):

So our \(x\) limits are between \(1\) from \(x=1\) and \(3\).

OpenStudy (anonymous):

Yup, i'm on board so far

OpenStudy (anonymous):

I suppose we should do shell method

OpenStudy (anonymous):

\[ 2\pi\int r(x)h(x)\;dx \]In this case \(r(x)=x\) and \(h(x)=3/x-x^2/9\)

OpenStudy (anonymous):

Does that make sense to you?

OpenStudy (anonymous):

Yeah, but when i'm solving for the Y-axis doesn't the x=1 play a role in the equation?

OpenStudy (anonymous):

Well

OpenStudy (anonymous):

and , if you don't mind me asking, is there certain scenarios where shell method is preferred over say, the washer method

OpenStudy (anonymous):